Jim covered 344 kilometers at the speed of 40 meters per hour. How long did Jim travel for in hours and in minutes?
Answer by CubeyThePenguin(3113)

speed = 40 m/hr = 0.04 km/hr

time taken = 344/0.04 = 8600 hr = 516000 min